OpenStudy (anonymous):

a) Write the next three terms and the twelfth term: 3, 12, 48… b) Describe the rule.

Parth (parthkohli):

Seems like we are multiplying 4 again and again, so this is a G.P.

Parth (parthkohli):

\(T_n = T_1 \times r^{n - 1} \Longrightarrow T_n = 3 \times 4^{n - 1}\)

Parth (parthkohli):

Now, you substitute n with 4,5,6 and 12.

OpenStudy (anonymous):

For 12th term use: \[\large a_{12} = (a) \cdot (r)^{n-1}\] Here a = 3, n = 12 and r = 4 Now find the 12th term..

Parth (parthkohli):

@waterineyes \(A_n\)*

OpenStudy (anonymous):

I told you you can write any symbol there..

OpenStudy (anonymous):

thanks guys

Parth (parthkohli):

You're welcome :)

OpenStudy (anonymous):

it does work but can u pls tell me how u work out the formula?